HUBBLE SPACE TELESCOPE DAILY REPORT #2136 PERIOD COVERED: 0000Z (UTC) 06/02/98 - 0000Z (UTC) 06/03/98 Daily Status Report as of 154/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Eight Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.2 Completed Seven Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.3 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 7618 (WF/PC-2 Cycle 7 Photometric Monitor UV/OPT STD) The WF/PC-2 was used to make observations of the standard star GRW+70D582 as part of a Cycle 7 photometric monitoring program. This ultraviolet spectrophotometric standard star was observed in a variety of filters and cameras to monitor the photometric stability and quantum efficiency of WF/PC-2 from the far-ultraviolet to near-infrared. The observations were completed as scheduled, and no problems were reported. 1.6 Completed Two Sets of Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.7 Completed NIC/1/2/3 7952 (Are Free-Floating Brown Dwarfs Single?) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the brown dwarf star DENISJ1228. Very recently the first two free-floating brown dwarfs have been discovered. High resolution Keck spectroscopy has confirmed strong lithium lines in both of them. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 7590 (Cosmological Parameters Omega and Lambda From High-Redshift Type IA Supernovae) The WF/PC-2 was used to make observations of type 1A supernova 14H0.9Z1 in order to study its redshift. The HST was used to make direct measurements of the cosmological parameters Omega, Lambda, and curvature using Type 1A supernovae as a calibrated "standard candle". The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.10 Completed NIC/1/2/3 7889 (Study of Compact Binaries in the Extreme Globular Cluster Terzan5) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of compact binary stars in the globular cluster TERZAN5. Stellar collisions and interactions in dense star clusters are fundamental to the evolution of globular clusters and galactic nuclei. The observations were completed as planned, and no anomalies were reported. 1.11 Completed Two Sets of STIS/CCD 7911 (STIS Galactic Plane Pure Parallel Program) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude survey observations in parallel with the prime science observations. Short exposures with the 50CCD imaging mode will provide the wavelength zero-point calibration. These data will be placed immediately into the Hubble Data Archive. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed FGS 7491 (The Masses of the Brown Dwarf Candidate GL 623B) The FGSs were used to conduct astrometry at 1-mass accuracy of faint companions to nearby M-dwarf star W922. This is the most productive direct method to determine accurate masses of stars near the end of the main sequence. From preliminary FGS results we estimate the mass of GL 623B as 0.10 solar masses, placing both near the stellar/substellar border. Such a mass determination for GL 623B should establish conclusively its nature as star or substellar object. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed NIC/1/2 7864 (The Host Galaxies of Radio-Quiet High- Redshift Quasars) The Near Infrared Camera (Cameras One and Two) was used to make observations of the radio-quite, high red shifted quasar MZZ1558 and its host galaxy. The underlying causes of the rapid cosmological evolution of the quasar population are a mystery around which are woven many different strands of contemporary cosmology. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Operations Notes: As documented in HSTAR 6502, the STIS MCE-2 reset at 153/190319Z while the low voltage was on and while in a level 4 SAA. ROP NS-12 was executed at 153/201433Z to clear the flight software error counter. The MCE-2 will be recovered via the SMS at the time of the next MAMA-2 turn-on (155/070013Z). Using ROP SR-1, the SSR EDAC error counter was reset two times. Several small segments of science SSR data were re-dumped for PACOR due to PSN errors in the original transmission. There were no reported problems in the re-dump. 3.0 SIGNIFICANT FORTHCOMING EVENTS: A GRGT-HST on-orbit test is scheduled from 154/1135Z until 154/2023Z. The second NICMOS Camera Three Campaign begins late tonight with the secondary mirror move scheduled to begin at approximately 155/0047Z. /DMH